# Voting-Based Semi-Parallel Proof-of-Work Protocol

### v1

@@ I Introduction
Nakamoto consensus [1], a solution to the decentralized consensus problem, has a vulnerability against withholding attacks such as double-spending [2] and selfish mining [3] due to the fact that the arrival of the blocks in the longest chain protocol are subject to Proof-of-Work (PoW) scheme which has an exponential distribution [4]. Since an exponential distribution with mean $\frac{1}{\lambda}$ has variance $\frac{1}{\lambda^{2}}$, adversaries can take advantage of the high variance of the block inter-arrival times by deviating from the honest behavior. More specifically, whenever it takes longer than usual for honest miners to mine a new block, the adversary can make profits from this by creating blocks and keeping them private until some predetermined conditions on the longest public chain are met [2, 3, 5].
@@
Many protocols have been suggested to solve some of the above problems of Nakamoto consensus. By modifying how transactions are appended to the ledger, Fruitchain [14] aims to increase the throughput level and achieve a more fair distribution of transaction rewards among miners with less variance. Bitcoin-NG [15] increases the throughput and reduces the variance of appending transactions to the ledger with a PoW leader election mechanism. GHOST rule [16] allows miners of the orphan blocks to get partial rewards by replacing the longest chain rule and mildens the unfair reward distribution. Modifications to fork choice rule of Nakamoto consensus such as [3, 17, 18, 19] try to mitigate the incentive attacks on reward mechanism. PHANTOM [20] and SPECTRE [21] protocols allow blocks to have multiple parents creating Directed Acyclic Graphs (DAG). As a result, higher block generation rates, $\lambda$, are allowed which in turn creates a more fair blockchain environment with high throughput. Multi-stage PoW [22] divides the PoW problem for each block into multiple sequential stages which decreases the variance of block interarrival times. Many of these protocols also come with limitations such as low fault tolerance, absence of rigorous security violation analysis, absence of an efficient, global and consistent ordering of transactions, communication overheads and resilience against possible withholding attacks [20, 21, 23, 24, 25, 26].
@@
Parallel PoW [27] has a similar principle where the system as a whole needs to solve $L$ puzzles in parallel, which decreases the variance of reward distribution times and increases resilience against double-spending by decreasing the variance of block interarrival times. However, in its raw form, parallel PoW protocols are more vulnerable to incentive attacks [28, 29]. Bobtail [4] can be seen as a modification to the raw form of parallel PoW scheme to increase the resilience against proof withholding attacks. Tailstorm and DAG-style voting schemes are other modifications introduced in [28, 29] that change the reward distribution rule between the parallel proofs to punish withholding attacks in the raw form of the parallel PoW scheme.
@@
In this work, we first consider the existing variations of the parallel PoW protocols [4, 28, 29] and develop hard-coded incentive attacks that show shortcomings of the protocols. For Bobtail [4], drawing an analogy to the analysis of Zhang and Preneel [19] about smallest hash based fork choice rules, we consider two types of attacks and argue that the the profitability threshold, $\alpha_{T}$, i.e., the minimum fraction of hashpower needed for a profitable attack, is zero. For Tailstorm [28], our hard-coded proof witholding attack results in a profitability threshold of $\alpha_{T}\leq\frac{1}{L}$ and results in revenues as large as the upper bound of the profit of a selfish miner in Nakamoto consensus. For DAG-style voting [29], even when the adversary has no network influence, i.e., $\gamma=0$, contrary to the authors’ intention, when $\alpha>\frac{1}{3}$, the punishment mechanism favors the adversary for a hard-coded attack we develop and the resilience is worse than the resilience in Tailstorm in this parameter regime. For all three schemes, we verify our theoretical analysis and arguments with simulations.
@@ III Proposed Protocol Description
Notice that, if a protocol separates the proofs from the ledgers, it decreases propagation times of proofs. As a result, less proofs will be mined during delay periods and thus the level of parallelness does not need to be high. This however, requires a ledger election mechanism that is incentive compatible. Further, in this case, ledger conflicts can be omitted with a clever mechanism as well. In this section, we explain a protocol that takes the superior properties of the existing PoW protocols such as [1, 20, 15, 18, 4, 27, 29]. The resulting protocol allows a reasonable amount of parallelness with small proof propagation times, an incentive compatible leader election mechanism, less communication overhead and has no ledger conflicts. Further, it has the superior variance properties compared to Nakamoto consensus, i.e., a more fair distribution of coins, better safety guarantees and resilience against withholding attacks and throughput increase.
@@ III-A Blocks and Chain Structure
Figure 3: The proof model of the semi-parallel PoW protocol.
@@ III-B Ledger Election
From $\omega^{h}_{i,4}$, it should be clear that within the cluster, each miner has a proposed ledger, and the ledger pays $\omega^{h}_{i,3}$ in transaction fees in addition to the coinbase reward. If we only accept one ledger, we need each proof of a cluster at height $h$ to agree with each other about the ledger of the previous height, in order to prevent rewriting of the history, i.e., double-spending. Thus, we consider the following scenario: The transaction fees associated with $\omega^{h}_{i,4}$ are offered to be shared with the miners of the proofs for the next block height $h+1$. If this is the case, it is in the best interest of miners at block height $h+1$ to pick the ledger of the previous height that pays the most fees.
Thus, the rule is to simply pick the ledger that pays the most among the valid ledgers of a cluster from the previous height, i.e., $\omega^{h}_{i,1}=\omega^{h-1}_{j^{*},4}$ where $j^{*}=\arg\max_{j}\{v(\omega^{h-1}_{j,3})\}$. Here, $v(\omega^{h-1}_{j,3}))=0$ if $\omega^{h-1}_{j,4}$ is not valid or withheld, i.e., the ledger is not shared with the honest miners in appropriate time, else $v$ is simply the identity function.
If parallel initiator proofs exist in the same cluster at block height $h$, i.e., $\omega^{h}_{1,2}=\omega^{h}_{2,2}$, then we require them to agree about their ledger pick for the previous height as well, i.e., $\omega^{h}_{1,1}=\omega^{h}_{2,1}$. As all incrementals build on top of the initiators, they necessarily inherit the pick of the initiators, hence we have a consensus about the ledger of the previous block height among the proofs clustered together. In Figure 3, the proofs that offer the most fees are represented with blue circles which are pointed by their creator as well as the initiator proof of the next block height. Furthermore, notice that, both initiator proofs at block $h$ pick the ledger of the proof $\omega_{5}^{h-1}$, hence they can be combined. On the other hand, even though the initiator proofs at block $h+1$ agree about the proofs in the previous block $h$, they pick different ledgers from the previous block, thus they cannot be combined.
@@ III-E Ledger Content
III-F Difficulty Adjustment and Fork Choice
@@
The miners pick the chain that contains the most aggregate work (heaviest chain). The aggregate work of a proof is the total number of proofs, weighted by their respective nonce difficulties it has in its history, plus the work of a compatible parallel proof if it exists. Here, without loss of generality, we assume that all the weights are the same. For example, the aggregate work of an incremental proof $\omega^{h}_{i}$ is the total number of proofs until block height $h$ plus the incremental height $\eta^{h}_{i,1}$ (plus one if there is another compatible proof $\omega^{h}_{j}$ with same history and $C(\omega^{h}_{i})=C(\omega^{h}_{j})$). We consider $b$-block confirmation rule where a ledger and the transactions inside are confirmed if the ledger is part of the heaviest chain and there are $b$ complete blocks mined on top of it.
@@ III-G Reward Distribution
Note that each proof in a cluster has the same difficulty contribution to the chain despite some might get half the reward for it (a punishment for parallelness). Thus, in total, the block at height $h$ receives mining rewards proportional to the total proof height, $C(\omega^{h}_{L})$, which is distributed between the miners of the block at height $h$. As an example, the block at height $h$ in Figure 3 receives $4$ units of rewards and each proof in this block gets $\frac{1}{2}$ units of rewards except the proof with incremental height $3$, which is the sole proof at proof height $2$ and gets $1$ unit of rewards. On the other hand, the block at height $h+2$ gets $7$ units of rewards and each proof gets $1$ unit of rewards.
@@ III-H Adversarial Ledgers
After the aggregate work at a cluster of block height h1h-1 reaches $L$ votes, the ledger of the proof that offers the most fees is expected to be shared with the network within a reasonable time (grace period) such as maximum network delays $\Delta+\delta$. Here, $\Delta$ is the maximum network delay when sending a transaction ledger of a block, whereas $\delta$ is the maximum network delay of sending a proof.
If the ledger that offers the most fees is not shared within the grace period, the ledger offering second most fees (and the others if necessary) is shared. At this point, honest miners can accept any valid ledger that has the most aggregate work and keep mining on it. Moreover, in this case, the proof associated with the ledger that was not shared can still be included as a valid vote and its aggregate work is counted towards the heaviest chain however the prover will be punished by not receiving the coinbase reward. Thus, it is in the interest of any miner (including the adversary) to share its ledger within the grace period if it offers the most fees.
In practice, the coinbase rewards are much higher than transaction fees thus it is not worth risking the coinbase rewards by creating false promises about the ledger fees. To decrease the effect of adversarial ledger not being shared in a grace period which slightly delays the transition to the next height, we could allow a bit more than $L+1$ proofs, where the reward split for extra proofs could be omitted as well if the ledger offering most fees is not shared. We leave these minor adjustments to specific protocol implementations.
@@ IV Analysis of the proposed protocol
In this section, we analyze the the superior properties our protocol inherits from the existing PoW protocols [1, 20, 15, 18]. We start analyzing the resilience against double-spending, i.e., the consistency. Then, we consider the proof withholding attacks and show the resilience using an MDP analysis. We also consider the fairness of the transaction fee distribution rule of our protocol.
@@ IV-A Consistency attacks
Here, we only analyze the double-spending probability for parameters $L\geq 10$, $b=1$ and give an upper bound for it. To repeat, our $b$-block confirmation rule states the following: A ledger together with the transactions inside are confirmed if it is the part of the heaviest chain and there are $b$-complete blocks mined on top of it. Notice that, in Nakamoto consensus, a $b$-block confirmation rule also counts the block that contains the ledger, therefore, when we compare our results, we should take this difference into account. More formally, our $b=1$ rule corresponds to $1\leq b\leq 2$ in Nakamoto consensus, since the ledger is proposed in one block and gets voted in the next (even though it remains hidden in the former block).
Essentially, we have to analyze the following attack: At block $h+1$, the miners start mining (voting) on ledgers that are created by miners of the proofs at block $h$. When an honest ledger of block $h$ receives $L$ votes at block $h+1$ and is part of the heaviest chain, the ledger is confirmed. See for example $\omega_{6}^{h}$ in Figure 3, which is confirmed when block $h+1$ is fully created. Any time after the honest ledger is confirmed, an adversarial ledger from block $h$, for example $\omega_{7}^{h}$ in Figure 3, is confirmed and replaces the honest ledger, if a heavier chain is published. Such a replacement is an instance of double-spending in our analysis.
Lemma 1
We skip proving this theorem as it essentially follows the same steps as bounded $\Delta$-delay model, by simply replacing blocks with proofs and $\Delta$ with $\delta$ in [30, Lemma 5]. We note that, since our protocol allows two parallel proofs to be combined, the balanced attacks are less likely to succeed and the rigged model assumption we make here is an extremely pessimistic one. Further, [7, 8, 9] show that more than two proofs being mined within a $\delta$ time is extremely unlikely, hence adversarial balanced attacks are less effective in our protocol.
$\lambda_{P}=L\lambda_{B},$ (1)
and
$p=\beta e^{-\lambda_{P}\delta}>\frac{1}{2}.$ (2)
Lemma 2
$S_{L}+S_{\delta_{L}}+S_{\Delta_{2}}.$ (3)
Lemma 3
The adversarial lead at $\tau$ can be upper bounded as
$M_{lead}\sim Geo\Big{(}1-\frac{q}{p}\Big{)}.$ (4)
Proof:  We refer the reader to [30, Section 5.1] for the details of the proof. Essentially, we can assume $\tau$ to be large and the lead becomes the steady-state distribution birth-death process since it cannot go below zero, where the probability of a step forward is 1p1-p, i.e., an adversarial arrival.  $\blacksquare$
Theorem 1
$P(M_{lead}+S_{L}+S_{\delta_{L}}+S_{\Delta_{2}}+M_{deficit}\geq L).$ (5)

### v2

@@ I Introduction
Nakamoto consensus [1], a solution to the decentralized consensus problem, has a vulnerability against withholding attacks such as double-spending [2] and selfish mining [3] due to the fact that the arrival of the blocks in the longest chain protocol are subject to Proof-of-Work (PoW) scheme which has an exponential distribution [4]. Since an exponential distribution with mean $\frac{1}{\lambda}$ has variance $\frac{1}{\lambda^{2}}$, adversaries can take advantage of the high variance of the block interarrival times by deviating from the honest behavior. More specifically, whenever it takes longer than usual for honest miners to mine a new block, the adversary can make profits from this by creating blocks and keeping them private until some predetermined conditions on the longest public chain are met [2, 3, 5].
@@
Many protocols have been suggested to solve some of the above problems of Nakamoto consensus. By modifying how transactions are appended to the ledger, Fruitchain [14] aims to increase the throughput level and achieve a more fair distribution of transaction rewards among miners with less variance. Sharding techniques such as Prism [Prism] try to increase throughput and reduce scalability and maintenance costs of miners. Bitcoin-NG [15] increases the throughput and reduces the variance of appending transactions to the ledger with a PoW leader election mechanism. GHOST rule [16] allows miners of the orphan blocks to get partial rewards by replacing the longest chain rule and mildens the unfair reward distribution. Modifications to fork choice rule of Nakamoto consensus such as [3, 17, 18, 19] try to mitigate the incentive attacks on reward mechanism. PHANTOM [20] and SPECTRE [21] protocols allow blocks to have multiple parents creating Directed Acyclic Graphs (DAG). As a result, higher block generation rates, $\lambda$, are allowed which in turn creates a more fair blockchain environment with high throughput. Multistage PoW [22] divides the PoW problem for each block into multiple sequential stages which decreases the variance of block interarrival times. Many of these protocols also come with limitations such as low fault tolerance, absence of rigorous security violation analysis, absence of an efficient, global and consistent ordering of transactions, communication overheads and resilience against possible withholding attacks [20, 21, 23, 24, 25, 26].
@@
Parallel PoW [27] has a similar principle where the system as a whole needs to solve $L$ puzzles in parallel, which decreases the variance of reward distribution times and increases resilience against double-spending by decreasing the variance of block interarrival times. However, in its raw form, parallel PoW protocols are more vulnerable to incentive attacks [28, 29]. Bobtail [4] can be seen as a modification to the raw form of parallel PoW scheme to increase the resilience against proof withholding attacks. Tailstorm and DAG-style voting schemes are other modifications introduced in [28, 29] that change the reward distribution rule between the parallel proofs to punish withholding attacks in the raw form of the parallel PoW scheme. It is worth noting that, parallel PoW protocols where miners create parallel proofs on top of the same block are distinct from sharding techniques where miners create and maintain parallel chains. In this paper, we only consider parallel PoW protocols.
@@
We first consider the existing variations of the parallel PoW protocols [4, 28, 29] and develop hard-coded incentive attacks that show shortcomings of the protocols. For Bobtail [4], drawing an analogy to the analysis of Zhang and Preneel [19] about smallest hash based fork choice rules, we consider two types of attacks and argue that the profitability threshold, $\alpha_{T}$, i.e., the minimum fraction of hashpower needed for a profitable attack, is zero. For Tailstorm [28], our hard-coded proof withholding attack results in a profitability threshold of $\alpha_{T}\leq\frac{1}{L}$ and results in revenues as large as the upper bound of the profit of a selfish miner in Nakamoto consensus. For DAG-style voting [29], even when the adversary has no network influence, i.e., $\gamma=0$, contrary to the authors’ intention, when adversarial fraction of hashpower $\alpha>\frac{1}{3}$, the punishment mechanism favors the adversary for a hard-coded attack we develop and the resilience is worse than the resilience in Tailstorm in this parameter regime. For all three schemes, we verify our theoretical analysis and arguments with simulations.
@@ III Proposed Protocol Description
Notice that, if a protocol separates the proofs from the ledgers, it decreases propagation times of proofs. As a result, less proofs will be mined during delay periods and thus the level of parallelness does not need to be high. This however, requires a ledger election mechanism that is incentive compatible. Further, in this case, ledger conflicts can be omitted with a clever mechanism as well. In this section, we explain a protocol that allows a reasonable amount of parallelness with small proof propagation times, an incentive compatible leader election mechanism, less communication overhead and has no ledger conflicts. Further, it has the superior variance properties compared to Nakamoto consensus, i.e., a more fair distribution of coins, better safety guarantees and resilience against withholding attacks and throughput increase.
@@ III-A Blocks and Chain Structure
Figure 3: The proof model of the semi-parallel PoW protocol with $L=7$. Orange and black circles represent the initiator and incremental proofs and the numbers inside the circles represent the incremental heights. Each initiator proof refers to a ledger from the previous block height as the winning ledger. Blue boxes represent the winning ledgers based on the most aggregate work fork choice rule. Here, if the proof $7$ of block $h+2$ is the most recent proof with most aggregate work and $b=1$, the ledgers until block height $h+1$ are confirmed, since they are part of the most aggregate work chain. But, ledger $\omega_{6}^{h+2}$ is not $b=1$-confirmed since 11-confirmation requires 11-complete block on height $h+3$ where the initiator of block $h+3$ needs to refer to $\omega_{6}^{h+2}$ as the winning ledger. Red dotted boxes (e.g., block $h$) together with a valid ledger (e.g., $\omega_{6}^{h}$) are considered as a complete block. Note that, block $h$ with ledger $\omega_{7}^{h}$ is a complete but orphan block. To replace the block $h$ and its associated ledger $\omega_{6}^{h}$ with the block $h$ and its associated ledger $\omega_{7}^{h}$ in the view of honest miners, the adversary creates an initiator proof that refers to $\omega_{7}^{h}$. The adversary has to create further proofs and share complete blocks at block height $h+1$ and $h+2$ to make the block $h$ and its associated ledger $\omega_{7}^{h}$ as the fork choice rule (most aggregate work). Note that the adversary cannot use any proofs of block $h+1$ or $h+2$ in the figure since initiator proof of the block $h+1$ refers to $\omega_{6}^{h}$ as the winning ledger and all the subsequent proofs build on top of this initiator proof. Thus, to rewrite the history, adversary has to work alone once it falls behind in the race of most aggregate work.
III-A1 Example proof model
@@ III-B Ledger Election
From $\omega^{h}_{i,4}$, it should be clear that within the cluster, each miner has a proposed ledger, and the ledger pays $\omega^{h}_{i,3}$ in transaction fees in addition to the coinbase reward. If we only accept one ledger, we need each proof of a cluster at height $h$ to agree with each other about the ledger of the previous block height, in order to prevent rewriting of the history, i.e., double-spending. Thus, we consider the following scenario: The transaction fees associated with $\omega^{h}_{i,4}$ are offered to be shared with the miners of the proofs for the next block height $h+1$. If this is the case, it is in the best interest of miners at block height $h+1$ to pick the ledger of the previous height that pays the most fees.
Thus, the rule is to simply pick the ledger that pays the most among the valid ledgers of a cluster from the previous height, i.e., $\omega^{h}_{i,1}=\omega^{h-1}_{j^{*},4}$ where $j^{*}=\arg\max_{j}\{v(\omega^{h-1}_{j,3})\}$. Here, $v(\omega^{h-1}_{j,3})=0$ if $\omega^{h-1}_{j,4}$ is not valid or withheld, else $v$ is simply the identity function.
If parallel initiator proofs exist in the same cluster at block height $h$, i.e., $\omega^{h}_{1,2}=\omega^{h}_{2,2}$, then we require them to agree about their ledger pick for the previous height as well, i.e., $\omega^{h}_{1,1}=\omega^{h}_{2,1}$. As all incrementals build on top of the initiators, they necessarily inherit the pick of the initiators, hence we have a consensus about the ledger of the previous block height among the proofs clustered together.
III-B1 Example proof model
In Figure 3, the proofs that offer the most fees are represented with blue circles which are pointed by their creator as well as the initiator proof of the next block height. Furthermore, notice that, both initiator proofs at block $h$ pick the ledger of the proof $\omega_{5}^{h-1}$, hence they can be combined. On the other hand, even though the initiator proofs at block $h+1$ agree about the proofs in the previous block $h$, they pick different ledgers from the previous block, thus they cannot be combined. Further, any incremental proof and blocks mined on top of one of these incremental proofs cannot be used by the other incremental proof since they inherit the ledger choice of the incremental proof. Thus, to rewrite the history, the adversary has to create a separate proof tree with more aggregate work.
@@ III-E Ledger Content
III-F Difficulty Adjustment, Fork Choice and Block Confirmation Rules
@@
At all times, the miners pick the chain that contains the most aggregate work (heaviest chain). The aggregate work of a proof is the total number of proofs, weighted by their respective nonce difficulties it has in its history, plus the work of a compatible parallel proof if it exists. Here, without loss of generality, we assume that all the weights are the same. For example, the aggregate work of an incremental proof $\omega^{h}_{i}$ is the total number of proofs until block height $h$ plus the incremental height $\eta^{h}_{i,1}$ (plus one if there is another compatible proof $\omega^{h}_{j}$ with same history and $C(\omega^{h}_{i})=C(\omega^{h}_{j})$). We consider $b$-block confirmation rule where a ledger and the transactions inside are confirmed if the ledger is part of the heaviest chain and there are $b$ complete blocks mined on top of it.
Note that, as the fork choice rule is the chain that contains the most aggregate work at all times, this implies that the ledger offering the most fees is only a temporary fork choice rule. More specifically, when a block with $L$ proofs is created and the ledger offering the most fees is shared, honest miners try to mine an initiator proof for the next block height referring to this ledger. However, as soon as the next initiator proof is created, this proof contains the most aggregate work and is the fork choice rule no matter which ledger it refers to as long as that ledger is semantically valid. We consider such a situation in more detail in Section D and show that it does not create any splits in honest views.
@@ III-G Reward Distribution
Note that each proof in a cluster has the same difficulty contribution to the chain despite some might get half the reward for it (a punishment for parallelness). Thus, in total, the block at height $h$ receives mining rewards proportional to the total proof height, $C(\omega^{h}_{L})$, which is distributed between the miners of the block at height $h$.
III-G1 Example proof model
The block at height $h$ in Figure 3 receives $4$ units of rewards and each proof in this block gets $\frac{1}{2}$ units of rewards except the proof with incremental height $3$, which is the sole proof at proof height $2$ and gets $1$ unit of rewards. On the other hand, the block at height $h+2$ gets $7$ units of rewards and each proof gets $1$ unit of rewards.
@@ III-H Adversarial Ledgers
After the aggregate work at a cluster of block height h1h-1 reaches $L$ votes, the ledger of the proof that offers the most fees is expected to be shared with the network within a reasonable time (grace period) such as maximum network delays $\Delta+\delta$. Here, $\Delta$ is the maximum network delay when sending a transaction ledger of a block, whereas $\delta$ is the maximum network delay of sending a proof. If the ledger that offers the most fees is not shared within the grace period, the ledger offering second most fees (and the others if necessary) is shared. Moreover, in this case, the proof associated with the ledger that was not shared can still be included as a valid vote and its aggregate work is counted towards the heaviest chain.
To decrease the effect of adversarial ledger not being shared in a grace period which slightly delays the transition to the next height, we could allow a bit more than $L+1$ proofs, where the reward split for extra proofs could be omitted as well if the ledger offering most fees is not shared. We leave these minor adjustments to specific protocol implementations. Note that, as it is common in the literature and the practice, we assume honest miners echo the messages they receive, hence all honest miners eventually receive the adversarial ledger if one of them receives.
@@ IV Analysis of the proposed protocol
In this section, we analyze the superior properties our protocol inherits from the existing PoW protocols [1, 20, 15, 18]. We start analyzing the resilience against double-spending, i.e., the consistency. Then, we consider the proof withholding attacks and show the resilience using an MDP analysis. We also consider the fairness of the transaction fee distribution rule of our protocol.
@@ IV-A1 System Model
IV-A2 Confirmation Rule
Here, we only consider the double-spending probability for parameters $L\geq 10$, $b=1$ and give an upper bound for it. To repeat, our $b$-block confirmation rule states the following: A ledger together with the transactions inside are confirmed if it is the part of the heaviest chain and there are $b$-complete blocks mined on top of it. Notice that, in Nakamoto consensus, a $b$-block confirmation rule also counts the block that contains the ledger, therefore, when we compare our results, we should take this difference into account. More formally, our $b=1$ rule corresponds to $1\leq b\leq 2$ in Nakamoto consensus, since the ledger is proposed in one block and gets voted in the next (even though it remains hidden in the former block).
We refer the reader to Appendix E for the description of the double-spending attack and the related analysis which relies on the state-of-the-art techniques for Nakamoto consensus [39, 30, 31, 40, 41].
@@ IV-A3 Numerical Evaluation
The results indicate a considerably superior security guarantees compared to Nakamoto consensus if we assume the same block sizes, generation times and network delays. More specifically, with $1$-block confirmation rule and $10$ minute block interarrival time, under $L=50$ and $\alpha=0.25$, an upper bound of double-spending probability around $10^{-4}$ is reached within less than $2$ blocks time in our protocol. In comparison, Bitcoin achieves a double-spending probability around $10^{-3}$ in $22$ blocks time[31].